If two like charges, each having a charge of four micro coulombs sit .6meters apart. The force between them will be?
1
Expert's answer
2020-05-20T09:31:54-0400

The Coulomb's force can be calculated the following ways:


"F=\frac{1}{4\pi \epsilon_0}\frac{q_1q_2}{r^2}=
\\\space\\
\frac{1}{4\pi 8.85\cdot10^{-12}}\frac{(4\cdot10^{-6})(4\cdot10^{-6})}{0.6^2}=0.4\text{ N}."